# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ Abstract
The cascade phenomenon is fundamental to turbulence, a typical non-equilibrium system. In turbulence, energy is cascaded from large to small scales, where it is dissipated by fluid viscosity. Using Illustris and Virgo simulations, this paper presents the cascade in another non-equilibrium system, the self-gravitating collisionless dark matter flow. The cosmic energy of dark matter decreases with time, as if "dissipated" due to the expanding background. This is facilitated by the energy cascade from large to small haloes via halo merging and from large to small scales in individual haloes via particle migration in fluctuating non-uniform gravitational potential, such that the cosmic energy is "dissipated" on small scales. We identify an inverse mass cascade across haloes of different sizes, which leads to a random walk of haloes in halo mass space. The halo mass function (double-$\lambda$) is naturally given by the corresponding Fokker-Planck equation for halo random walk. Similarly, the random walk of particles in haloes leads to the distribution of particles. The halo density profile (double-$\gamma$) can be analytically derived from the corresponding Fokker-Planck equation for particle random walk. Universal scaling laws were identified that exhibit small-scale permanence for the halo density. The different inner density slopes of simulated haloes can be explained by the nonzero net mass and energy flux in individual haloes. The mass and energy cascade in dark matter flow establishes a statistically steady state to continuously release the system energy and maximize the system entropy. The key feature of this statistically steady state is scale-independent rates of cascade such that the statistical structures of the haloes are self-similar and scale-free, and there is no net accumulation of mass and energy on any intermediate scales. Since the waiting time and jumping length for particle random walk depends on the particle mass $m_{p}$, new mass constraints can be identified as $m_{p}\geq 10^{-15}$kg or $10^{12}$eV. This constraint excludes the standard WIMPs and suggests a heavy dark matter scenario (superheavy right-handed neutrinos, etc.).
@@
Dark matter; N-body simulations; Theoretical models
@@ 1 Introduction
Collisionless systems often have properties suggesting common physical principles that control their motion and evolution. The self-gravitating collisionless fluid dynamics (SG-CFD) is the study of these principles for the motion of collisionless matter under the influence of gravity. Typical examples can be the large-scale gravitational collapse of collisionless dark matter, which is a non-equilibrium self-gravitating collisionless flow problem (Lukic et al., 2007). Gravitational instability leads to the self-organization of collisionless dark matter particles and the formation and evolution of large-scale structures. Within a CDM (cold dark matter) paradigm (Blumenthal et al., 1984), the initial density fluctuation has a larger amplitude at smaller scales. The formation of structures starts from the gravitational collapse of small-scale density fluctuations and proceeds hierarchically such that small structures coalesce into large structures in a "bottom-up" fashion. Highly localized, over-dense, and virialized haloes are major manifestations of the nonlinear gravitational collapse (Neyman and Scott, 1952; Cooray and Sheth, 2002) and the building blocks of large-scale structures, whose abundance and internal structures have been extensively studied over the last several decades.
@@
The abundance of haloes is often described by a halo mass function, one of the most fundamental quantities to probe large-scale structures and model the structure formation and evolution. The first landmark might be the Press-Schechter (PS) formalism (Press and Schechter, 1974; Bond et al., 1991), which allows one to predict the shape and evolution of mass function. The distribution of halo mass is determined by postulating that the probability of forming haloes is related to the amplitude of density fluctuations on that scale. Haloes will form at some mass scale once the smoothed linear density contrast on that scale exceeds a threshold value $\delta_{c}$. This threshold value must be analytically derived by examining the nonlinear collapse of a spherical top hat over-density (Tomita, 1969; Gunn and Gott, 1972). Alternative derivations using an excursion set approach (EPS) put the PS formalism on a firmer footing by removing the fudge factor in the original PS model (Bond et al., 1991). This was further extended to the excursion set with correlated steps (Musso and Sheth, 2012; Paranjape and Sheth, 2012; Maggiore and Riotto, 2010). Although mathematically less rigorous, PS formalism and its extensions are still very useful and allow one to compute many different structural properties. Examples are the halo mass function, merging rates, and clustering properties.
@@
The halo internal structures are usually characterized by the halo density profile (Del Popolo and Kroupa, 2009), which can be analyzed both analytically and through numerical N-body simulations (Moore et al., 1998; Klypin et al., 2001). Since the first study on the spherical collapse (Gunn and Gott, 1972), the power-law density profile was proposed using the self-similar approximation (Bertschinger, 1985; Fillmore and Goldreich, 1984). High-resolution N-body simulations have revealed a nearly universal profile exhibiting a cuspy density—less steep than the isothermal profile at smaller radii and becomes steeper at larger radii (Navarro et al., 1997, 2004). However, there has yet to be a consensus on the exact slope value of the inner density slope $s$ obtained via N-body simulations. Since the first slope $s=-1.0$ described by the NFW profile (Navarro et al., 1997), inner slopes of the simulated haloes have varied in a wide range, including values such as $s>-1.0$ (Navarro et al., 2010), $s=-1.2$ (Diemand and Moore, 2011), and $s\approx-1.3$ (Governato et al., 2010; McKeown et al., 2022; Lazar et al., 2020). When compared with observations, the predicted cuspy inner density tend to be steeper (de Blok and Bosma, 2002; de Blok et al., 2003; Swaters et al., 2003; Kuzio de Naray and Kaufmann, 2011). This gives rise to the cusp-core problem, one of the so-called small scale challenges.
@@
To the author’s knowledge, the exact origin of the nearly universal density profile and the different inner density slopes of simulated haloes are not yet fully understood (Cooray and Sheth, 2002). This paper treats the halo mass function and density profiles from a different perspective. By revisiting the cascade in turbulence, we reveal the existence of a mass and energy cascade in the collisionless dark matter flow. The mass cascade describes the mass transfer between haloes on different mass scales, which leads to the distribution of haloes with respect to the halo mass, i.e., the halo mass function. The energy cascade in haloes describes the energy flow along the halo radial direction, which suggests the distribution of particles in haloes, i.e., the halo density profile. This new perspective offers a theory for the near-universal halo mass functions and density profiles, as well as the widespread inner density slopes of simulated haloes.
@@
Turbulence is a typical non-equilibrium system. Constant chaotic fluid motion and energy flowing through different scales prevent turbulence from reaching a stable equilibrium state. The energy cascade is fundamental to turbulence. At high Reynolds numbers, turbulence consists of a random collection of eddies (building blocks) at different length scales that interact with each other and dynamically change, which can be described by a famous poem (Richardson, 1922):
@@
The poem describes a conceptual picture that large eddies feed smaller eddies, which feed even smaller eddies and then lead to viscous dissipation at the smallest scale, i.e., the concept of a direct energy cascade. There is a broad spectrum of eddy sizes within fully developed turbulence. Large eddies are usually created by the instability of large-scale mean flow at integral scale $L$. They rapidly break up and pass their kinetic energy to smaller eddies due to inertial force. Smaller eddies are transient and, in turn, pass their energy to even smaller eddies. The cascade continues down the scale and stops to operate at the smallest eddies (dissipation scale $\eta$), where the viscous force becomes dominant over the inertial force. At high Reynolds numbers, there exists a range of length scales where the viscous force is negligible and the inertial force is dominant. The rate $\varepsilon$ of energy passing down the cascade should be scale-independent in this range and match exactly the rate of energy dissipation at the smallest scale. The direct energy cascade is a dominant feature of three-dimensional turbulence. However, an inverse energy cascade was predicted for two-dimensional turbulence, where kinetic energy is transferred from small to large scales (Kraichnan, 1967).
From this brief description, turbulence has two key features enabling the cascade phenomenon: i) A broad spectrum of eddies that mediate the energy cascade across different scales; ii) The viscous force operating on small scales to dissipate the system energy into heat (radiation) and maintain a steady energy cascade at a constant rate. Apparently, as another example of the non-equilibrium system, the self-gravitating collisionless dark matter flow (SG-CFD) shares some striking similarities with turbulence.
First, there also exists a broad spectrum of halo sizes to maximize the system entropy (Xu, 2023a). Small haloes are created by gravitational instability and interacting and merging with other haloes. Haloes pass their mass onto larger and larger haloes. Consequently, we expect a continuous cascade of mass from small to large mass scales. By simply replacing "whirls" in that poem with "haloes," a new poem now describes the inverse mass cascade from small to large scales that is consistent with the hierarchical structure formation:
@@
Second, the total energy of the collisionless dark matter in an expanding background constantly decreases with time, as shown by the cosmic energy evolution (Section 3). Since there is no viscous force operating in collisionless dark matter, the decrease in total energy is caused by the Hubble expansion (Eq. (6)). In this regard, the Hubble parameter $H$ plays a "similar" role as the viscosity in turbulence that leads to the energy "dissipation" in collisionless dark matter flow. The quotation indicates the "dissipation" here is different.

### v4

# Universal spectrum and scaling laws for halo mass function and structureC.6

@@ Abstract
[…]
@@
Dark matter; N-body simulations; Theoretical models; Universal spectrum; Scaling laws;
@@ 1 Introduction
Collisionless systems often have properties that suggest common physical principles governing their motion and evolution. The self-gravitating collisionless fluid dynamics (SG-CFD) is the study of these principles for the motion of collisionless matter under the influence of gravity. A typical example is the gravitational collapse of collisionless dark matter, a non-equilibrium self-gravitating collisionless flow problem (Lukic:2007-The-halo-mass-function--High-r). Gravitational instability leads to the self-organization of collisionless dark matter particles and the structure formation and evolution. Within a CDM paradigm (cold dark matter), the initial density fluctuation has a larger amplitude at smaller scales (Blumenthal:1984-Formation-of-Galaxies-and-Larg). The formation of structures starts from the gravitational collapse of small-scale density fluctuations and proceeds hierarchically so that small structures coalesce into large structures in a "bottom-up" fashion.
The hierarchical structure formation has been studied in great detail since the 1980s using cosmological N-body simulations of increasing resolution and fidelity. The vast majority of these simulations adopt Gaussian initial conditions with a linear power spectrum of cold, warm, or hot dark matter variants. Evolution from these initial conditions is governed by the coupled collisionless Boltzmann (Vlasov) and Poisson equations. Consequently, any statistical properties are primarily consequences of the Gaussian statistics together with the shape of the initial linear power spectrum. This is reflected in most analytic and/or semi-analytical approaches, such as Press–Schechter formalisms and merger trees, where the abundance and internal structures of halos are direct reflections of the initial spectrum and cosmological conditions.
In the linear regime, gravity drives the linear amplification of density perturbations. The growth rate of perturbations is scale-independent, i.e., each mode grows independently and proportionally to the linear spectrum as $\propto a^{2}$, where $a$ is the scale factor. Consequently, the abundance of the halo and the internal structure retain quantitative memory of the initial distribution of fluctuations. From this perspective, structure formation reflects a continuous interplay between gravitational dynamics and the statistical imprint of the initial power spectrum: gravity governs the dynamics of collapse, while the initial spectrum determines the sequence and characteristic scales at which the collapse unfolds.
In the nonlinear regime, because of the "bottom-up" nature, smaller haloes form earlier under conditions that are more directly reflective of the initial spectrum. Massive haloes are formed later through the continuous hierarchical merging and smooth accretion of matter. This prolonged and violent assembly history, which includes many mergers and phases of violent relaxation, acts to "wash out" the imprints of the initial spectrum. Therefore, small scales retain more memory of the shape of the initial spectrum than large scales. Merging enabled mass and energy redistribution (or "cascade") across different scales progressively erases the imprint of the primordial spectrum. In this paper, we thereby propose that with increasing scale or mass of haloes, gravitational dynamics ultimately dominate, establishing a quasi-universal range where the shape of the spectrum is fully determined by gravity and no longer depends on the initial conditions. The enabling process, a "cascade" of mass and energy across different scales, is also proposed and formulated.
To better demonstrate the idea, Figure 1 presents the evolution of the density power spectrum $P(k,z)$ with redshifts z from the Virgo SCDM simulations (Frenk:2000-Public-Release-of-N-body-simul). The linear power spectrum $P_{L}(k)$ is plotted as a dotted line. Three distinct regimes and two critical scales can be identified: 1) a linear regime on scales $k<k_{NL1}$, where scale $k_{NL1}$ is determined by the linear dimensionless power spectrum $\Delta_{L}^{2}(k_{NL1})=1$, at which fluctuations are just turning nonlinear; 2) a universal transition range (thick solid lines) on scales $k_{NL1}<k<k_{NL2}$ with a spectrum index $n\approx$-1, where gravitational dynamics dominate over the effects of the initial spectrum. The critical scale $k_{NL2}$ is determined by extending the spectrum nn=-1 (dashed line) from $k_{NL1}$ to intersect with the actual spectrum $P(k,z)$ at $k_{NL2}$; 3) a fully nonlinear range on scales $k>k_{NL2}$ that retains the full memory of the shape of the initial spectrum. The universal range with nn=-1 is gradually formed at a high redshift and expands with time, with an increasing ratio $k_{NL2}/k_{NL1}\propto a^{1/2}$. The scale $k_{h}^{*}\approx$ 1h/Mpc corresponds to the size of the halo of the characteristic mass $m^{*}_{h}$ at z=0, which is located in the universal transition range with $k_{NL1}<k_{h}^{*}<k_{NL2}$. The haloes of mass $m_{h}^{*}$ rapidly collapse, merge, and virialize, with the most intense energy and mass transfer. This is the most gravitationally violent scale that is primarily dominated by gravitational dynamics, leading to a universal spectrum index nn=-1 around this scale.
Figure 1: The density power spectrum $P(k,z)$ with comoving wavenumber $k$ at different redshifts z from Virgo SCDM simulations. The dotted line denotes the linear power spectrum $P_{L}(k)$. The figure presents three distinct regimes and two critical scales: 1) a linear regime on scales $k<k_{NL1}$, where scale $k_{NL1}$ denotes the deviation from the linear theory; 2) a universal transition range (thick solid lines) on scales $k_{NL1}<k<k_{NL2}$ with $P(k)\propto k^{-1}$, where the gravitational dynamics dominate over the the initial spectrum; 3) a fully nonlinear range on scales $k>k_{NL2}$ that retains the memory of the initial spectrum. While both scales decrease with time, the ratio $k_{NL2}/k_{NL1}$ increases with time, reflecting the expanding universal range. The scale $k_{h}^{*}\approx$ 1h/Mpc corresponds to the size of haloes of characteristic mass $m^{*}_{h}$ at z=0, such that these haloes are primarily dominated by the gravitational dynamics, not the initial conditions. This paper focuses on the universal range and associated mass and energy cascade across scales, and their effects on structure formation and evolution.
The existence of such a universal and expanding transition range is a direct result of the mass and energy cascade across scales. As the mass and energy are transferred across scales through many nonlinear couplings, the detailed structure of the initial spectrum becomes statistically irrelevant. Therefore, scales in the universal range no longer “remember” the shape of the initial spectrum, leading to universal abundance and structure for haloes near $m_{h}^{*}$ in this range. Thus, the loss of memory of the initial spectrum is a direct consequence of the mass/energy cascade. Therefore, we also aim to develop a cascade theory leading to this quasi-universal range, which is distinct from the Press–Schechter formalisms. To facilitate the discussion, we start with a brief overview of the current state of knowledge.
@@
Highly localized, over-dense, and virialized haloes are major manifestations of the nonlinear gravitational collapse (Neyman:1952-A-Theory-of-the-Spatial-Distri; Cooray:2002-Halo-models-of-large-scale-str) and the building blocks of large-scale structures, whose abundance and internal structures have been extensively studied over the last several decades. The abundance of haloes is often described by a halo mass function, one of the most fundamental quantities to probe large-scale structures and model the formation and evolution of the structure. The first landmark might be the Press-Schechter (PS) formalism (Press:1974-Formation-of-Galaxies-and-Clus; Bond:1991-Excursion-Set-Mass-Functions-f). The distribution of halo mass is determined by postulating that the probability of forming haloes is related to the amplitude of density fluctuations on that scale. Haloes will form on some mass scale once the smoothed linear density contrast on that scale exceeds a threshold value. This threshold value must be analytically derived by examining the nonlinear collapse of a spherical top hat over-density (Tomita:1969-Formation-of-Gravitationally-B; Gunn:1972-Infall-of-Matter-into-Clusters). Alternative derivations using an excursion set approach (EPS) put the PS formalism on a firmer footing by removing the fudge factor in the original PS model (Bond:1991-Excursion-Set-Mass-Functions-f). This was further extended to the excursion set with correlated steps (Musso:2012-One-step-beyond-the-excursion-set-approach; Paranjape:2012-Peaks-theory-and-the-excursion-set-approach; Maggiore:2010-THE-HALO-MASS-FUNCTION-FROM-EXCURSION-SET-THEORY). Although mathematically less rigorous, PS formalism and its extensions are still very useful and allow one to compute many different structural properties. Examples are the halo mass function, merging rates, and clustering properties.
@@
The internal structures of the halo are generally characterized by the halo density profile (Del_Popolo:2009-Density-profiles-of-dark-matte), which can be analyzed both analytically and numerically through N-body simulations (Moore:1998-Resolving-the-structure-of-col; Klypin:2001-Resolving-the-structure-of-col). Since the first study on spherical collapse (Gunn:1972-Infall-of-Matter-into-Clusters), the power-law density profile was proposed using the self-similar approximation (Bertschinger:1985-Self-Similar-Secondary-Infall-; Fillmore:1984-Self-Similar-Gravitational-Col). High-resolution N-body simulations have revealed a nearly universal profile exhibiting a cuspy density—less steep than the isothermal profile at smaller radii and becomes steeper at larger radii (Navarro:1997-A-universal-density-profile-fr; Navarro:2004-The-inner-structure-of-ACDM-ha). However, there has yet to be a consensus on the exact slope value of the inner density slope $s$ obtained by N-body simulations. Since the first slope s=1.0s=-1.0 described by the NFW profile (Navarro:1997-A-universal-density-profile-fr), inner slopes of the simulated haloes have varied in a wide range, including values such as s>1.0s>-1.0 (Navarro:2010-The-diversity-and-similarity-of-simulated), s=1.2s=-1.2 (Diemand:2011-The-Structure-and-Evolution-of-Cold-Dark), and $s\approx-1.3$ (Governato:2010-Bulgeless-dwarf-galaxies-and-dark-matter-cores; McKeown:2022-Amplified-J-factors-in-the-Galactic-Centre; Lazar:2020-A-dark-matter-profile-to-model-diverse). Compared with observations, the predicted inner density of the cuspy tends to be higher (Blok:2002GALAXIES:-STRUCTURE-GALAXIES; Blok:2003-Simulating-observations-of-dark-matter; Swaters:2002-The-Central-mass-distribution-in-dwarf; Naray:2011-Recovering-cores-and-cusps-in-dark-matter; Salucci:2019-The-distribution-of-dark-matter). This gives rise to the cusp-core problem, one of the so-called small scale challenges.
@@
To our knowledge, the exact origin of the nearly universal density profile and the different inner density slopes of the simulated haloes is not yet fully understood (Cooray:2002-Halo-models-of-large-scale-str). This paper treats the halo mass function and density profiles from a different perspective. We first establish the existence of a universal range and mass and energy cascade in the collisionless dark matter flow. The mass cascade describes the mass transfer between haloes across different mass scales, which leads to the distribution of haloes with respect to the halo mass, i.e., the halo mass function. The energy cascade in haloes describes the energy flow along the halo radial direction, which suggests the distribution of particles, i.e., the halo density profile. This new perspective offers a theory for nearly universal halo mass functions and density profiles. To illustrate and compare, we begin by revisiting the cascade phenomenon in turbulence.
@@
Turbulence is a typical non-equilibrium system, a long-standing and probably the last unresolved problem in classical physics. Constant chaotic fluid motion and energy flowing through different scales prevent the turbulence from reaching a stable equilibrium state. The energy cascade is fundamental to turbulence. At high Reynolds numbers, turbulence consists of a random collection of eddies (building blocks) at different length scales that interact with each other and dynamically change, which can be described by a famous poem (Richardson:1922-Weather-Prediction-by-Numerica):
@@
The poem describes a conceptual picture that large eddies feed smaller eddies, which feed even smaller eddies and then lead to viscous dissipation at the smallest scale, i.e., the concept of a direct energy cascade. There is a broad spectrum of eddy sizes within fully developed turbulence. Large eddies are usually created by the instability of large-scale mean flow at an integral scale $L$. They rapidly break up and pass their kinetic energy to smaller eddies as a result of the inertial force. Smaller eddies are transient and, in turn, pass their energy to even smaller eddies. The cascade continues down the scale and stops operating in the smallest eddies (dissipation scale $\eta$), where the viscous force becomes dominant over the inertial force. At high Reynolds numbers, there exists a range of length scales in which the viscous force is negligible and the inertial force is dominant. The rate $\varepsilon$ of energy passing through the cascade should be scale-independent in this range and match exactly the rate of energy dissipation at the smallest scale. The direct energy cascade is a dominant feature of three-dimensional turbulence. However, an inverse energy cascade was predicted for two-dimensional turbulence, where kinetic energy is transferred from small to large scales (Kraichnan:1967-Inertial-Ranges-in-2-Dimension).
From this brief description, turbulence has two key features that enable the cascade phenomenon: i) a broad spectrum of eddies that mediate the energy cascade across different scales; and ii) a viscous force operating on small scales to dissipate the system energy into heat (radiation) and maintain a steady energy cascade at a constant rate. Through energy cascade and dissipation, the turbulence, as a typical nonequilibrium system, reaches a statistically steady state to continuously release system energy and maximize entropy. This intermediate statistical equilibrium state is manifested by a scale-independent rate of the energy cascade $\varepsilon$.
Apparently, the self-gravitating collisionless dark matter flow (SG-CFD) is different from fluid turbulence. Although SG-CFD is incompressible in full six-dimensional phase space, its three-dimensional fluid projection lacks the dissipative mechanisms and the vortex mechanism to enable the same energy cascade as turbulence. Therefore, SG-CFD cannot directly admit an eddy-mediated energy cascade in a viscous fluid. However, collisionless dark matter flow also shares striking similarities with turbulence, allowing a different, halo-mediated mass and energy "cascade". Note that the term “cascade” is used in a generalized sense to denote the redistribution of mass and energy across scales, rather than a turbulent viscous process.
First, there also exists a broad spectrum of halo sizes to maximize the entropy of the system (Xu:2023-Maximum-entropy-distributions-of-dark-matter). Small haloes are created by gravitational instability and interacting and merging with other haloes. Haloes pass their mass onto larger and larger haloes. Consequently, we expect a continuous cascade of mass from small to large scales. By simply replacing "whirls" in that poem with "haloes," a new poem now describes the inverse mass cascade from small to large scales, which is consistent with hierarchical structure formation:
@@
Second, based on Noether’s theorem, energy is not globally conserved in an expanding universe because constantly stretching space is not symmetric in time. The total energy of collisionless dark matter constantly decreases with time during gravitational collapse in an expanding background, as shown by the cosmic energy evolution (Section B). Since there is no viscous force operating in collisionless dark matter, the decrease in total energy is caused by the Hubble expansion (Eq. (121)). In this regard, the Hubble parameter $H$ plays a "similar" role as the viscosity in turbulence that leads to the energy "dissipation" in collisionless dark matter flow, which initiates and sustains a continuous energy cascade across different scales. However, unlike turbulence, the reader should note that the term "dissipation" is used here to stand for the energy decrease. No energy is truly dissipated into radiation during structure formation.